order: acetaminophen suspension 170mg po prn q4h for tempature >102

available :300 mg /5ml
how many ml will you give ?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

           Formula = Desired dose/Dose in hand * Quantity.

     Here

          Desired dose(D) = 170 mg.

          Dose in hand (H) = 300 mg.

              Quantity(Q) = 5 ml.

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 170/300 * 5

                                                     = 0.56 * 5

                                                     = 2.8 ml.